# Coupled dimerized alternating-bond quantum spin chains in the distorted honeycomb-lattice magnet Cu5SbO6

### v1

@@ Appendix A Dimer expansion for triplet excitations
$\boldsymbol{a}_{2}={}$ $l_{2}(\cos\phi\boldsymbol{e}_{x}+\sin\phi\boldsymbol{e}_{y}),$ (6)
$\boldsymbol{a}_{3}={}$ $l_{3}(\cos\theta\boldsymbol{e}_{x}+\sin\theta\boldsymbol{e}_{z}).$ (7)
@@
$\boldsymbol{b}_{1}={}$ $\frac{2\pi}{l_{1}}(\boldsymbol{e}_{x}-\cot\phi\boldsymbol{e}_{y}-\cot\theta\boldsymbol{e}_{z}),$ (8)
$\boldsymbol{b}_{2}={}$ $\frac{2\pi}{l_{2}}\csc\phi\boldsymbol{e}_{y},$ (9)
$\boldsymbol{b}_{3}={}$ $\frac{2\pi}{l_{3}}\csc\theta\boldsymbol{e}_{z}.$ (10)
$\boldsymbol{b}_{1}$ and $\boldsymbol{b}_{2}$ form the reciprocal dimer lattice of the $J_{1}$-$J_{2}$-$J_{3}$ model whereas $\boldsymbol{b}_{1}$ and $\boldsymbol{b}_{3}$ form that of the $J_{1}$-$J_{2}$-$J_{4}$ model.
@@ A.1 Energy dispersion
$-J_{3}\cos\left(\frac{\boldsymbol{q}\cdot\boldsymbol{a}_{1}}{2}\right)\cos\left[\boldsymbol{q}\cdot\left(\boldsymbol{a}_{2}-\frac{1}{2}\boldsymbol{a}_{1}\right)\right].$
Defining $\boldsymbol{q}\equiv\left(2\pi q_{h}/l_{1}\right)\boldsymbol{e}_{x}+(\pi q_{k}\csc\phi/l_{2})\boldsymbol{e}_{y}$ leads to
$\boldsymbol{q}\cdot\boldsymbol{a}_{1}=2\pi q_{h},\quad\boldsymbol{q}\cdot\left(\boldsymbol{a}_{2}-\frac{1}{2}\boldsymbol{a}_{1}\right)=\pi q_{k},$
where we set $\phi=\pi/6$ and $l_{1}=\sqrt{3}l_{2}$. Therefore,
$\Delta E_{\boldsymbol{q}}^{123}=J_{1}{}$ $-\frac{J_{2}}{2}\cos(2\pi q_{h})$
$-J_{3}\cos\left(\pi q_{h}\right)\cos\left(\pi q_{k}\right).$ (11)
Our result for $\Delta E_{\boldsymbol{q}}^{123}$ agrees with the expansion of Eq. (1) in Ref. [Gao2020] with respect to $J_{2}$ and $J_{3}$. Similarly, for the $J_{1}$-$J_{2}$-$J_{4}$ model,
$\Delta E_{\boldsymbol{q}}^{124}=J_{1}-\frac{J_{2}}{2}\cos(\boldsymbol{q}\cdot\boldsymbol{a}_{1})-\frac{J_{4}}{2}\cos[\boldsymbol{q}\cdot(\boldsymbol{a}_{3}-\boldsymbol{a}_{1})].$
Defining $\boldsymbol{q}\equiv\left(2\pi q_{h}/l_{1}\right)\boldsymbol{e}_{x}+(2\pi q_{l}\csc\theta/l_{3})\boldsymbol{e}_{z}$ leads to
$\boldsymbol{q}\cdot\boldsymbol{a}_{1}=2\pi q_{h},\quad\boldsymbol{q}\cdot\left(\boldsymbol{a}_{3}-\boldsymbol{a}_{1}\right)=2\pi q_{l}-2\pi q_{h}(1-\xi),$
where $\xi\equiv(l_{3}/l_{1})\cos\theta$. Therefore,
$\Delta E_{\boldsymbol{q}}^{124}=J_{1}{}$ $-\frac{J_{2}}{2}\cos(2\pi q_{h})-\frac{J_{4}}{2}\cos\{2\pi[q_{l}-(1-\xi)q_{h}]\}.$
Although $\Delta E_{\boldsymbol{q}}^{124}$ includes a parameter $\xi$ related to the lattice geometry, it can be eliminated from the expression of the density of states. By performing a variable transformation $q_{l}=q_{l}^{\prime}+(1-\xi)q_{h}$, we have
$\Delta\tilde{E}^{124}(q_{h},q_{l}^{\prime})\equiv{}$ $\Delta E^{124}(q_{h},q_{l}^{\prime}+(1-\xi)q_{h})$
$={}$ $J_{1}-\frac{J_{2}}{2}\cos(2\pi q_{h})-\frac{J_{4}}{2}\cos(2\pi q_{l}^{\prime}).$ (12)
Therefore, the density of states can be written as
$D(\omega)={}$ $\int_{0}^{1}dq_{h}\int_{0}^{1}dq_{l}\;\delta(\omega-\Delta E(q_{h},q_{l}))$
$={}$ $\int_{0}^{1}dq_{h}\int_{-(1-\xi)q_{h}}^{1-(1-\xi)q_{h}}dq_{l}^{\prime}\;\delta(\omega-\Delta\tilde{E}(q_{h},q_{l}^{\prime}))$
$={}$ $\int_{0}^{1}dq_{h}\left[\int_{-(1-\xi)q_{h}}^{0}dq_{l}^{\prime}\;\delta(\omega-\Delta\tilde{E}(q_{h},q_{l}^{\prime}))+\int^{1-(1-\xi)q_{h}}_{0}dq_{l}^{\prime}\;\delta(\omega-\Delta\tilde{E}(q_{h},q_{l}^{\prime}))\right]$
$={}$ $\int_{0}^{1}dq_{h}\left[\int_{1-(1-\xi)q_{h}}^{1}dq_{l}^{\prime\prime}\;\delta(\omega-\Delta\tilde{E}(q_{h},q_{l}^{\prime\prime}-1))+\int^{1-(1-\xi)q_{h}}_{0}dq_{l}^{\prime}\;\delta(\omega-\Delta\tilde{E}(q_{h},q_{l}^{\prime}))\right]$
$={}$ $\int_{0}^{1}dq_{h}\int_{0}^{1}dq_{l}^{\prime}\;\delta(\omega-\Delta\tilde{E}(q_{h},q_{l}^{\prime})).$
@@ A.2 Dynamical structure factor
$\boldsymbol{q}\cdot(\boldsymbol{d}_{1}-\boldsymbol{d}_{0})=\frac{4\pi}{3}q_{h},$
@@
$|\langle t^{+}_{\boldsymbol{q}}|S^{+}_{\boldsymbol{q}}|s\rangle|^{2}=1-\cos\left(\frac{4\pi}{3}q_{h}\right).$
@@
$\mathcal{S}^{+-}(\boldsymbol{q},\omega)=\left[1-\cos\left(\frac{4\pi}{3}q_{h}\right)\right]\delta(\omega-\Delta E_{\boldsymbol{q}}).$

### v2

@@ Appendix A Dimer expansion for triplet excitations
$\boldsymbol{a}_{2}={}$ $l_{2}(\cos\phi_{\mathrm{cso}}\boldsymbol{e}_{x}+\sin\phi_{\mathrm{cso}}\boldsymbol{e}_{y}),$ (6)
$\boldsymbol{a}_{3}={}$ $l_{3}(\cos\theta_{\mathrm{cso}}\boldsymbol{e}_{x}+\sin\theta_{\mathrm{cso}}\boldsymbol{e}_{z}).$ (7)
@@
$\boldsymbol{b}_{1}={}$ $\frac{2\pi}{l_{1}}(\boldsymbol{e}_{x}-\cot\phi_{\mathrm{cso}}\boldsymbol{e}_{y}-\cot\theta_{\mathrm{cso}}\boldsymbol{e}_{z}),$ (8)
$\boldsymbol{b}_{2}={}$ $\frac{2\pi}{l_{2}}\csc\phi_{\mathrm{cso}}\boldsymbol{e}_{y},$ (9)
$\boldsymbol{b}_{3}={}$ $\frac{2\pi}{l_{3}}\csc\theta_{\mathrm{cso}}\boldsymbol{e}_{z}.$ (10)
$\boldsymbol{b}_{1}$ and $\boldsymbol{b}_{2}$ form the reciprocal dimer lattice of the $J_{1}$-$J_{2}$-$J_{3}$ model whereas $\boldsymbol{b}_{1}$ and $\boldsymbol{b}_{3}$ form that of the $J_{1}$-$J_{2}$-$J_{4}$ model. Note that for Cu5SbO6, $l_{1}=9.00$ Å, $l_{2}=5.30$ Å, $l_{3}=6.55$ Å, $\theta_{\mathrm{cso}}=62^{\circ}$, $\phi_{\mathrm{cso}}=30^{\circ}$.
@@ A.1 Energy dispersion
$-J_{3}\cos\left(\frac{\boldsymbol{Q}\cdot\boldsymbol{a}_{1}}{2}\right)\cos\left[\boldsymbol{Q}\cdot\left(\boldsymbol{a}_{2}-\frac{1}{2}\boldsymbol{a}_{1}\right)\right].$ (11)
Similarly, for the $J_{1}$-$J_{2}$-$J_{4}$ model,
$\Delta E_{\boldsymbol{Q}}^{124}=J_{1}-\frac{J_{2}}{2}\cos(\boldsymbol{Q}\cdot\boldsymbol{a}_{1})-\frac{J_{4}}{2}\cos[\boldsymbol{Q}\cdot(\boldsymbol{a}_{3}-\boldsymbol{a}_{1})].$ (12)
Defining $\boldsymbol{Q}=q_{1}\boldsymbol{b}_{1}+q_{2}\boldsymbol{b}_{2}+q_{3}\boldsymbol{b}_{3}$, we then have
$\boldsymbol{Q}\cdot\boldsymbol{a}_{j}=\sum_{i}q_{i}(\boldsymbol{b}_{i}\cdot\boldsymbol{a}_{j})=\sum_{i}q_{i}(2\pi\delta_{ij})=2\pi q_{j}.$ (13)
Eqs. (11) and (12) become
$\Delta E^{\mathrm{123}}_{\boldsymbol{Q}}={}$ $J_{1}-\frac{J_{2}}{2}\cos(2\pi q_{1})-J_{3}\cos(\pi q_{1})\cos[\pi(2q_{2}-q_{1})],$ (14)
$\Delta E^{\mathrm{124}}_{\boldsymbol{Q}}={}$ $J_{1}-\frac{J_{2}}{2}\cos(2\pi q_{1})-\frac{J_{4}}{2}\cos[2\pi(q_{3}-q_{1})].$ (15)
Figure 8: (a)-(d) Dispersions along $q_{1}$ and $q_{2}$ directions for the $J_{1}$-$J_{2}$-$J_{3}$ model calculated with $J_{1}=16.95$ meV, $J_{2}=-3.9$ meV, and $J_{3}=0.9$ meV. (e)-(h) Dispersions along $q_{1}$ and $q_{3}$ directions for the $J_{1}$-$J_{2}$-$J_{4}$ model calculated with $J_{1}=16.5$ meV, $J_{2}=-6$ meV, and $J_{4}=3$ meV.
@@ A.2 Dynamical structure factor
$\boldsymbol{Q}\cdot(\boldsymbol{d}_{1}-\boldsymbol{d}_{0})=\frac{4\pi}{3}q_{1},$
@@
$|\langle t^{+}_{\boldsymbol{Q}}|S^{+}_{\boldsymbol{Q}}|s\rangle|^{2}=1-\cos\left(\frac{4\pi}{3}q_{1}\right).$
@@
$\mathcal{S}^{+-}(\boldsymbol{Q},\omega)=\left[1-\cos\left(\frac{4\pi}{3}q_{1}\right)\right]\delta(\omega-\Delta E_{\boldsymbol{Q}}).$ (18)